Solve by the elimination method.
0.3x - 0.2y = 4
0.5x + 0.3y = 1
What is the solution of the system?
(Type and ordered pair.)
(Type an integer or a fraction. Type N if there is no solution. Type I if there are infinitely many solutions.)

Answer by checkley77(12844)

0.3x - 0.2y = 4 multiply by 3.
0.5x + 0.3y = 1 muyltiply by 2 & then add them.
.9x-.6y=12
1x+.6y=2
-----------------------------
1.9x=14
x=14/1.9
x=7.368 Answer.
.3*7.368-.2y=4
2.21-.2y=4
-.2y=4-2.21
-.2y=1.79
y=1.79/-.2
y=-8.95 Answer.
Proof:
.5*7.368+.3*-8.95=1
3.685-2.685=1
1=1